OpenStudy (anonymous):

2/3(x-4)=2x I would only like some guidance on this please!

OpenStudy (anonymous):

distribute 2/3 to x-4

OpenStudy (anonymous):

2/3x - 8/3 or 2.666666

OpenStudy (anonymous):

8/3 would be easiest to solve for x

OpenStudy (anonymous):

okay so it would be 2/3x-8/3=2x add 8/3 on both sides?

OpenStudy (anonymous):

no. you want to get x on one side

OpenStudy (anonymous):

so divide 2/3x on both sides

OpenStudy (anonymous):

which would be -8/3= 2x/ (2/3x) but since you can not divide 2/3 you want to times 2x by the reciprocal of 2/3 which would be....?

OpenStudy (anonymous):

you can not divide 2x by 2/3x*

OpenStudy (anonymous):

3/2...so okay i got the rest now, the answer is -2, am i right?

OpenStudy (anonymous):

i got -8

OpenStudy (anonymous):

-8/3=2x/1*3/2x -8/3=3x -8/3*3/1=x -8=x

OpenStudy (anonymous):

oh, i accidentally wrote down the wrong fraction. But i understand now, thank you!! (: